# Attention as an Adaptive Filter

### v1

@@ 3.7 Unitary Dynamics
$\lim_{\mathrm{Re}(\lambda_{k})\to 0}\frac{1-e^{2\mathrm{Re}(\lambda_{k})\Delta t\tau}}{-2\mathrm{Re}(\lambda_{k})}=\Delta t\,\tau$
@@
$\mathcal{K}^{V}(\tau)=\sigma_{V}^{2}(\tau)\mathbf{I}_{d},\quad\text{where:}\quad\sigma_{V}^{2}(\tau):=\lambda_{C}^{2}\sigma^{2}\Delta t\tau+\eta^{2}$
And the $ij$th component of the precision matrix becomes:
$\mathbf{P}_{ij}=\frac{1}{\sigma_{V}^{2}(|i-j|)},$
Since the eigenvalues are purely imaginary, $\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|=\|e^{\bm{\lambda}t_{i}}\|=1$, and: $\|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\|=\|\mathbf{Z}_{\mathbf{k}}[k,j]\|$, $\|\mathbf{\hat{Z}}_{\mathbf{q}}[k,j]\|=\|\mathbf{Z}_{\mathbf{q}}[k,j]\|$. Hence:
$\|\mathbf{R_{qk}}[i,j]\|^{2}=\sum_{k}\Big{(}\|\mathbf{Z}_{\mathbf{q}}[k,i]\|^{2}+\|\mathbf{Z}_{\mathbf{k}}[k,j]\|^{2}-2\mathrm{Re}\big{(}\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big{)}\Big{)}$
If, in addition, $\mathbf{S}$ is unitary (in which case $\mathbf{A}$ is skew-symmetric), then $\|\mathbf{Z}_{\mathbf{q}}[k,i]\|^{2}=\|\mathbf{Z}[k,i]\|^{2}$, $\|\mathbf{Z}_{\mathbf{v}}[k,i]\|^{2}=\|\mathbf{Z}[k,i]\|^{2}$, so if the inputs have been normalized to unit length, then: $\sum_{k}\|\mathbf{Z}_{\mathbf{q}}[k,i]\|^{2}=\sum_{k}\|\mathbf{Z}_{\mathbf{k}}[k,j]\|^{2}=1$. Hence,
$\|\mathbf{R_{qk}}\|^{2}=2\Big{(}1-\mathrm{Re}\big{(}\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big{)}\Big{)}$
@@
$\mathbf{L}=-\log\Big(\nu\mathbf{V}^{C}+2-2\mathrm{Re}\big{(}\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big{)}\Big{missing})$
@@
If we set $\sigma^{2}\rightarrow 0$ (zero process noise), then $\mathbf{V}^{C}=\eta^{2}\mathbf{1}$ becomes a simple scalar bias. Letting $\tau=2/(\nu\eta^{2}+2)$, the unnormalized softmax becomes:
@@
The constant term $\frac{\tau}{2}$ outside the exponential cancels in the normalization. The imaginary part of the matrix exponential becomes a multiplicative relative positional encoding, and we recover a complex-valued form of ordinary dot-product attention.
3.8 Implementation
3.8.1 Adaptive Filter Attention

### v2

@@ 3.8 Unitary Dynamics
$\lim_{\alpha\to 0}\frac{1-e^{2\alpha\Delta t\tau}}{-2\alpha}=\Delta t\,\tau$
@@
$\sigma_{V}^{2}(\tau):=\big|\lambda_{C}\big|^{2}\sigma^{2}\Delta t\tau+\eta^{2}$
If, in addition, $\mathbf{S}$ is unitary (in which case $\mathbf{A}$ is skew-symmetric and the dynamics are unitary in the original basis), then $\big|\mathbf{Z}_{\mathbf{q}}[k,i]\big|^{2}=\big|\mathbf{Z}[k,i]\big|^{2}$, $\big|\mathbf{Z}_{\mathbf{v}}[k,i]\big|^{2}=\big|\mathbf{Z}[k,i]\big|^{2}$, so if the inputs are normalized to unit length, the queries and keys are as well: $\sum_{k}\big|\mathbf{Z}_{\mathbf{q}}[k,i]\big|^{2}=\sum_{k}\big|\mathbf{Z}_{\mathbf{k}}[k,j]\big|^{2}=1$. Hence,
$\big|\mathbf{R_{qk}}\big|^{2}=2\Big(1-\mathrm{Re}\big(\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big)\Big)$
@@
[…]
@@
If we set $\sigma^{2}\rightarrow 0$ (zero process noise), i.e. assume a deterministic system, then $\mathbf{V}^{C}=\eta^{2}\mathbf{1}$ becomes a simple scalar bias. Letting $\tau=2/(\nu\eta^{2}+2)$, the unnormalized softmax becomes:
@@
The constant term $\frac{\tau}{2}$ outside the exponential cancels in the normalization. The imaginary part of the matrix exponential becomes a multiplicative relative positional encoding, of the form of xPos (Sun et al., 2022) and RetNet (Sun et al., 2023). Under the assumption $\alpha=0$, the propagated covariance $\mathbf{V}^{C}$ is a bias of the form a|ij|+ba|i-j|+b, as in ALiBi (Press et al., 2022).
In the above formulation, if the inputs are normalized beforehand, then the projection matrices $\mathbf{W_{q}}$ and $\mathbf{W_{k}}$ must be unitary to ensure that $|\mathbf{\hat{Z}}_{q}|$ and $|\mathbf{\hat{Z}}_{k}|$ remain 1, simplifying the computation. By contrast, normalizing the queries and keys after the projections removes this constraint, since $\mathbf{W_{q}}$ and $\mathbf{W_{k}}$ no longer need to preserve norms. This motivates the choice of post-projection normalization, a strategy explored in prior work such as Query-Key Normalization for Transformers (Henry et al., 2020).
3.9 The Inhomogeneous Case
Consider the inhomogeneous SDE:
$d\mathbf{x}(t)=\big(\mathbf{A}\mathbf{x}(t)+\mathbf{u}\big)\,dt+\mathbf{G}\,d\mathbf{w}(t),\qquad\mathbf{x},\mathbf{u}\in\mathbb{R}^{d};\,\mathbf{A},\mathbf{G}\in\mathbb{R}^{d\times d}$
where $\mathbf{u}$ is a constant drift term.
We can represent this as a homogeneous SDE with one additional dimension:
$d\mathbf{x}_{\mathrm{aug}}(t)=\mathbf{A}_{\mathrm{aug}}\,\mathbf{x}_{\mathrm{aug}}(t)\,dt+\mathbf{G}_{\mathrm{aug}}\,d\mathbf{w}(t),$
$\mathbf{x}_{\mathrm{aug}}(t)=\begin{bmatrix}\mathbf{x}(t)\\[2.0pt] 1\end{bmatrix}\in\mathbb{R}^{d+1},\quad\mathbf{A}_{\mathrm{aug}}=\begin{pmatrix}\mathbf{A}&\mathbf{u}\\[4.0pt] \mathbf{0}^{\top}&0\end{pmatrix},\quad\mathbf{G}_{\mathrm{aug}}=\begin{pmatrix}\mathbf{G}\\[4.0pt] \mathbf{0}^{\top}\end{pmatrix}\in\mathbb{R}^{(d+1)\times(d+1)}.$
However, $\mathbf{A}_{\mathrm{aug}}$ and $\mathbf{Q}_{\mathrm{aug}}=\mathbf{G}_{\mathrm{aug}}\mathbf{G}_{\mathrm{aug}}^{\top}$ are now singular. Hence, we must deal with the inhomogeneous case directly.
The solution to the above inhomogeneous SDE is:
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i})+\left(\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}(\Delta t_{ij}-\tau)}d\tau\right)\mathbf{u}+\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}(\Delta t_{ij}-\tau)}\mathbf{G}d\mathbf{w}(\tau)$
where $\Delta t=t_{j}-t_{i}$. Letting $\mathbf{G}_{u}(\Delta t_{ij}):=\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}(\Delta t_{ij}-\tau)}d\tau$, our pulled-back/forward estimate of $\mathbf{x}(t_{i})$ from $\mathbf{x}(t_{j})$ is then:
$\mathbf{\hat{x}}_{i}(t_{j})=e^{-\mathbf{A}\Delta t_{ij}}\big(\mathbf{x}(t_{j})-\mathbf{G}_{u}(\Delta t_{ij})\mathbf{u}\big)$
Letting $\mathbf{u_{s}}:=\mathbf{S}\mathbf{u}\mathbf{S}^{-1}$, the drift term is:
$\mathbf{G}_{u}(\Delta t_{ij})\mathbf{u}=\mathbf{S}\Big(e^{\bm{\Lambda}\Delta t_{ij}}\int_{0}^{\Delta t_{ij}}e^{-\bm{\Lambda}\tau}d\tau\Big)\mathbf{u}_{\mathbf{s}}\mathbf{S}^{-1}=\mathbf{S}\Big(\frac{e^{\bm{\Lambda}\Delta t_{ij}}-\mathbf{I}}{\bm{\Lambda}}\Big)\mathbf{u}_{\mathbf{s}}\mathbf{S}^{-1}$
In the eigenbasis,
$\mathbf{\hat{x}}_{\mathbf{s}i}(t_{j})=e^{-\bm{\Lambda}\Delta t_{ij}}\Big(\mathbf{x}_{\mathbf{s}}(t_{j})-\frac{e^{\bm{\Lambda}\Delta t_{ij}}-\mathbf{I}}{\bm{\Lambda}}\Big)\mathbf{u}_{\mathbf{s}}$
$e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{x}_{\mathbf{s}}(t_{j})-\Big(\frac{\mathbf{I}-e^{-\bm{\Lambda}\Delta t_{ij}}}{\bm{\Lambda}}\Big)\mathbf{u}_{\mathbf{s}}$
$e^{-\bm{\Lambda}\Delta t_{ij}}\Big(\mathbf{x}_{\mathbf{s}}(t_{j})+\frac{\mathbf{u}_{\mathbf{s}}}{\bm{\Lambda}}\Big)-\frac{\mathbf{u}_{\mathbf{s}}}{\bm{\Lambda}}$
(where, as always, division is element-wise).
Including the measurement model:
$\mathbf{\hat{z}}_{\mathbf{s}i}(t_{j})=e^{-\bm{\Lambda}\Delta t_{ij}}\Big(\mathbf{z}_{\mathbf{s}}(t_{j})+\bm{\Lambda}_{C}\frac{\mathbf{u}_{\mathbf{s}}}{\bm{\Lambda}}\Big)-\bm{\Lambda}_{C}\frac{\mathbf{u}_{\mathbf{s}}}{\bm{\Lambda}}$
Since $\mathbf{u}$ is deterministic, the propagation of uncertainty is unchanged.
We will now examine how this modifies AFA. Define a bias:
$\mathbf{b}_{\mathbf{qk}}[k]:=\bm{\lambda}_{C,k}\frac{\mathbf{u}_{\mathbf{qk}}[k]}{\bm{\lambda}_{\mathbf{qk},k}}$
Our key estimate tensor is:
$\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]=\mathcal{E}_{\mathbf{qk}}[k,i,j]\cdot\Big(\mathbf{Z_{k}}[k,j]+\mathbf{b}_{\mathbf{qk}}[k]\Big)-\mathbf{b}_{\mathbf{qk}}[k]$
So the residual tensor is:
$\Big(\mathbf{Z_{q}}[k,i]+\mathbf{b}_{\mathbf{qk}}[k]\Big)-\mathcal{E}_{\mathbf{qk}}[k,i,j]\cdot\Big(\mathbf{Z_{k}}[k,j]+\mathbf{b}_{\mathbf{qk}}[k]\Big)$
Define biased queries and keys:
$\mathbf{Z}^{u}_{\mathbf{q}}[k,j]:=\mathbf{Z_{q}}[k,i]+\mathbf{b}_{\mathbf{qk}}[k],\quad\mathbf{Z}^{u}_{\mathbf{k}}[k,j]:=\mathbf{Z_{k}}[k,j]+\mathbf{b}_{\mathbf{qk}}[k].$
The residual becomes:
$\mathcal{R}_{\mathbf{qk}}[k,i,j]=\mathbf{Z}^{u}_{\mathbf{q}}[k,j]-\mathcal{E}_{\mathbf{qk}}[k,i,j]\cdot\mathbf{Z}^{u}_{\mathbf{k}}[k,j]$
$\bm{\Phi}_{\mathbf{qk}}[k,i]\cdot\left(\mathbf{\hat{Z}}^{u}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}_{\mathbf{k}}^{u}[k,j]\right)$
We can then proceed exactly as before, using $\mathbf{\hat{Z}}^{u}_{\mathbf{q}}[k,i]$ and $\mathbf{\hat{Z}}_{\mathbf{k}}^{u}[k,j]$ in place of $\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]$ and $\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]$.
The value estimate tensor is:
$\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]=\mathcal{E}_{\mathbf{v}}[k,i,j]\cdot\Big(\mathbf{Z_{v}}[k,j]+\mathbf{b}_{\mathbf{v}}[k]\Big)-\mathbf{b}_{\mathbf{v}}[k]$
So the MLE is:
$\mathbf{\bar{Z}_{v}}[k,i]:=\sum_{j\leq i}\mathcal{A}[k,i,j]\cdot\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$
$\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\mathcal{A}[k,i,j]\cdot\mathbf{\hat{Z}}^{u}_{\mathbf{v}}[k,j]-\mathbf{b}_{\mathbf{v}}[k]\cdot\sum_{j\leq i}\mathcal{A}[k,i,j]$
$\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\mathbf{A}[i,j]\cdot\mathbf{\hat{Z}}^{u}_{\mathbf{v}}[k,j]-\mathbf{b}_{\mathbf{v}}[k]$
Finally, the prediction step becomes:
$\mathbf{Z}_{\mathbf{v}}^{+}[k,i]=e^{\bm{\lambda}_{\mathbf{v},k}\Delta t}\cdot\mathbf{\bar{Z}}[k,i]+\mathbf{b}_{\mathbf{v}}[k]\Big(e^{\bm{\lambda}_{\mathbf{v},k}\Delta t}-1\Big)$
$e^{\bm{\lambda}_{\mathbf{v},k}\Delta t}\cdot\Big(\mathbf{\bar{Z}}[k,i]+\mathbf{b}_{\mathbf{v}}[k]\Big)-\mathbf{b}_{\mathbf{v}}[k]$
Hence,
$\mathbf{Z}_{\mathbf{v}}^{+}[k,i]=e^{\bm{\lambda}_{\mathbf{v},k}\Delta t}\cdot\Big(\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\mathbf{A}[i,j]\cdot\mathbf{\hat{Z}}^{u}_{\mathbf{v}}[k,j]\Big)-\mathbf{b}_{\mathbf{v}}[k]$
Hence, we can handle the inhomogeneous case simply by including bias terms in the definitions of the queries, keys, values, and predictions:
$\mathbf{Z_{q}}=\mathbf{W_{q}}\mathbf{Z}+\mathbf{b_{q}},\quad\mathbf{Z_{k}}=\mathbf{W_{k}}\mathbf{Z}+\mathbf{b_{k}},\quad\mathbf{Z_{v}}=\mathbf{W_{v}}\mathbf{Z}+\mathbf{b_{v}},\quad\mathbf{Z}^{+}=\mathbf{W_{p}}\mathbf{Z}_{\mathbf{v}}^{+}+\mathbf{b_{p}}.$
3.10 Measurement Time Delay and Bias
Let us now consider the effect of a time delay $\phi_{t}$ and a bias term $\mathbf{b_{z}}$ in our measurement model:
$d\mathbf{x}(t)=\big(\mathbf{A}\mathbf{x}(t)+\mathbf{u}\big)dt+\mathbf{G}d\mathbf{w}(t),\quad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k}+\phi_{t})+\mathbf{b}_{\mathbf{z}}+\mathbf{v}(t_{k}),$
Our estimate of the state at time $t_{j}$ from the measurement at time $t_{j}$ is now:
$\mathbf{x}^{\text{est}}(t_{j}+\phi_{t})=\mathbf{C}^{-1}\big(\mathbf{z}(t_{j})-\mathbf{b}_{\mathbf{z}}\big)$
We can then estimate the state at time $t_{j}$ from the estimate of the state at time $t_{j}+\phi_{t}$:
$\mathbf{x}^{\text{est}}(t_{j})=e^{-\mathbf{A}\phi_{t}}\big(\mathbf{x}^{\text{est}}(t_{j}+\phi_{t})-\mathbf{G}_{u}(\phi_{t})\mathbf{u}\big)$
We can then estimate the state at time $t_{i}$ from the state at time $t_{j}$:
$\mathbf{x}^{\text{est}}(t_{i})=e^{\mathbf{-A}\Delta t_{ij}}\big(\mathbf{x}^{\text{est}}(t_{j})-\mathbf{G}_{u}(\Delta t_{ij})\mathbf{u}\big)$
We then estimate the state at time $t_{i}+\phi_{t}$ from the estimate of the state at time $t_{i}$:
$\mathbf{x}^{\text{est}}(t_{i})=e^{-\mathbf{A}\phi_{t}}\big(\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})-\mathbf{G}_{u}(\phi_{t})\mathbf{u}\big)$
Hence,
$\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})=e^{\mathbf{A}\phi_{t}}\mathbf{x}^{\text{est}}(t_{i})+\mathbf{G}_{u}(\phi_{t})\mathbf{u}$
Substituting the expression for $\mathbf{x}^{\text{est}}(t_{i})$:
$\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})=e^{\mathbf{A}\phi_{t}}\Big(e^{\mathbf{-A}\Delta t_{ij}}\big(\mathbf{x}^{\text{est}}(t_{j})-\mathbf{G}_{u}(\Delta t_{ij})\mathbf{u}\big)\Big)+\mathbf{G}_{u}(\phi_{t})\mathbf{u}$
$e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{x}^{\text{est}}(t_{j})+\Big(\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}$
Substituting the expression for $\mathbf{x}^{\text{est}}(t_{j})$:
$\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})=e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\Big(e^{-\mathbf{A}\phi_{t}}\big(\mathbf{x}^{\text{est}}(t_{j}+\phi_{t})-\mathbf{G}_{u}(\phi_{t})\mathbf{u}\big)\Big)+\Big(\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}$
$e^{\mathbf{-A}\Delta t_{ij}}\mathbf{x}^{\text{est}}(t_{j}+\phi_{t})+\Big(\big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\big)\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}$
Substituting the expression for $\mathbf{x}^{\text{est}}(t_{j}+\phi_{t})$:
$\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})=e^{\mathbf{-A}\Delta t_{ij}}\Big(\mathbf{C}^{-1}\big(\mathbf{z}(t_{j})-\mathbf{b}_{\mathbf{z}}\big)\Big)+\Big(\big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\big)\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}$
Finally, we estimate the measurement at time $t_{i}$ from the estimated state at time $t_{i}+\phi_{t}$:
$\mathbf{\hat{z}}(t_{i})=\mathbf{C}\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})+\mathbf{b}_{\mathbf{z}}$
$\mathbf{C}\Big[e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}\big(\mathbf{z}(t_{j})-\mathbf{b}_{\mathbf{z}}\big)+\Big(\big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\big)\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}\Big]+\mathbf{b}_{\mathbf{z}}$
$\underbrace{\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}\mathbf{z}(t_{j})}_{\text{{Propagated measurement}}}+\underbrace{\mathbf{C}\Big(\big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\big)\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}}_{\text{Propagated drift}}+\underbrace{\Big(\mathbf{I}-\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}\Big)\mathbf{b}_{\mathbf{z}}}_{\text{Propagated bias}}$
If $\mathbf{A}$ and $\mathbf{C}$ commute, this becomes:
$\mathbf{\hat{z}}(t_{i})=e^{\mathbf{-A}\Delta t_{ij}}\mathbf{z}(t_{j})+\mathbf{C}\Big(\big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\big)\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}+\Big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\Big)\mathbf{b}_{\mathbf{z}}$
Let’s examine the drift term (in the eigenbasis):
$\big(\mathbf{I}-e^{\bm{-\Lambda}\Delta t_{ij}}\big)\Big(\frac{e^{\bm{\Lambda}\phi_{t}}-\mathbf{I}}{\bm{\Lambda}}\Big)-e^{\bm{-\Lambda}(\Delta t_{ij}-\phi_{t})}\Big(\frac{e^{\bm{\Lambda}\Delta t_{ij}}-\mathbf{I}}{\bm{\Lambda}}\Big)$
$\Big(\frac{e^{\bm{\Lambda}\phi_{t}}-\mathbf{I}}{\bm{\Lambda}}\Big)-e^{\bm{-\Lambda}\Delta t_{ij}}\Big(\frac{e^{\bm{\Lambda}\phi_{t}}-\mathbf{I}}{\bm{\Lambda}}\Big)-e^{\bm{\Lambda}(\phi_{t})}\Big(\frac{\mathbf{I}-e^{\bm{-\Lambda}\Delta t_{ij}}}{\bm{\Lambda}}\Big)$
$\frac{\mathbf{I}-e^{\bm{-\Lambda}\Delta t_{ij}}}{\bm{\Lambda}}$
Hence,
$\mathbf{\hat{z}}_{\mathbf{s}}(t_{i})=e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{z}_{\mathbf{s}}(t_{j})+\Big(\mathbf{I}-e^{-\bm{\Lambda}\Delta t_{ij}}\Big)\Big(\mathbf{b}_{\mathbf{z,s}}-\bm{\Lambda}_{C}\frac{\mathbf{u}_{\mathbf{s}}}{\bm{\Lambda}}\Big)$
So the time delay term disappears and we just get an extra constant bias term, which can be absorbed into the query, key, value, and prediction biases, as before.
4 Implementation
4.1 Algorithm